Activity-Based Supplier Costing

Ventana Company is a car window repair and replacement company operating in the after-sales market. Ventana’s purchasing manager uses two suppliers (Jones Glass and Claro Glass) for the source of its passenger car windows. Data relating to side windows (Side) and windshields (WS) are given below.

I. Activity costs

Activity
Adverse buying* 648,000
Supplier returns** 71,500

* Extra cost of purchasing from local car dealer because of insufficient delivery of supplier.
** Windows returned because they were not ordered or because they were defective.

II. Supplier Data

Jones Glass Claro Glass
Side WS Side WS
Unit purchase price $80 $135 $77 $132
Units purchased 15,000 15,000 30,000 30,000
Insufficient units 750 750 3,300 3,300
Returned units 375 375 1,000 1,000

Required:

1. Calculate the activity rates for assigning costs to suppliers.

Adverse buying rate $ per adverse purchase
Supplier return rate $ per return

2. Calculate the total unit purchasing cost for each component for each supplier. If required, round your answers to the nearest cent.

Total unit cost
Jones Glass Claro Glass
Side $ $
WS $ $
Answer #1
Required 1 :
Total insufficient units = 750 + 750 + 3300 + 3300 8100
Total returned units = 375 + 375 + 1000 + 1000 2750
Adverse buying rate = Adverse buying cost / Total insufficient units = 648000 / 8100 80 per adverse purchase
Supplier return rate = Supplier return cost / Total returned units = 71500 / 2750 26 per return
Required 2 :
Allocation of activity cost :
Jones Glass Claro Glass
Side WS Side WS

Adverse buying cost

60000

[ 750*80 ]

60000

[ 750*80 ]

264000

[ 3300*80 ]

264000

[ 3300*80 ]

Supplier returns cost

9750

[ 375*26 ]

9750

[ 375*26 ]

26000

[ 1000*26 ]

26000

[ 1000*26 ]

Total activity cost 69750 69750 290000 290000
(/) Units purchased 15000 15000 30000 30000
Activity cost per unit 4.65 4.65 9.67 9.67

Total unit cost

[ Unit purchase price + Activity cost per unit ]

Jones Glass Claro Glass

Side

84.65

[ 80+4.65 ]

86.67

[ 77+9.67 ]

WS

139.65

[ 135+4.65 ]

141.67

[ 132+9.67 ]